If the area of a square with side x is 9, what is the area of a square of side 4x?
Answer by richard1234(7193)

The side length is multiplied by 4, so the area is multiplied by 4^2, or 16. The new area is 9*16 = 144.

Or we can solve for x, x^2 = 9, x = 3, (4x)^2 = 144.